1. A city block is 3 times as long as it is wide. If the distance around the block is 0.48 kilometers, what is the area of the block in square meters?
Answer by ankor@dixie-net.com(22740)

1. A city block is 3 times as long as it is wide.
If the distance around the block is 0.48 kilometers,
what is the area of the block in square meters?
:
Change .48 km to 480 meters
Let x = the width of the block
then
3x = the length
Find x
2(3x) + 2x = 480
simplify, divide by 2
3x + x = 240
x = 240/4
x = 60 m is the lenth
then
3(60) = 180 m is the length
Find the area
180 * 60 = 10,800 sq/m is the area